OpenStudy (meilendurcer):

how to solve 4/3 = 8/x

OpenStudy (tkhunny):

Have you considered multiplying by 3? See if that helps.

OpenStudy (compassionate):

Hello, you want to first cross multiply. That is: \[\frac{ 4 }{ 3 } = \frac{ 8 }{ x } \rightarrow \frac{ 4 \times x }{ 3 \times 8 } \]

OpenStudy (anonymous):